Q.2 For the nitrogen molecule, calculate the following: i) Translation energy per molecule at 300K ii) Translation energy per mole at 300K iii) Number density at 298.2K iv) Concentration at 298.2K
1
Expert's answer
2014-03-14T05:10:04-0400
KE per molecule = 3/2 kT = 3/2 * 1.38e-23 * 300 = 6.21e-21
(J) & KE per mole: 3/2 RT = 3.74 kJ In the next calculations, it is suggested that the pressure is equal to the
standard pressure. Let's suggest that the volume of the gas at 298.2 K is 1 m3. At 273.2 K the gas
would have a volume of 1*273.2/298.2 = 0.916 m3 which is 40.9 moles = 2.46e25
molecules. So, The ND is 2.46e25 molecules / m3 at 298.2K The concentration is 40.9 moles / m3